OpenStudy (anonymous):

Line CD contains points A (4, −7) and B (4, 8). The slope of line CD is zero undefined positive negative

OpenStudy (math&ing001):

According to slope formula: slope=8-(-7)/4-4=15/0 which is undefined.

OpenStudy (anonymous):

ok thank u @math&ing001

OpenStudy (math&ing001):

Welcome !